# On the Robustness of Language Models for Tabular Question Answering

### v2

@@ Appendix G Additional Robustness Result
The table 4 summarizes the impact of various perturbation operations on the accuracy and robustness of larger and smaller LLMs models when performing TQA across three distinct datasets (WTQ, TAT-QA, and SCITAB). For each perturbation categories the EM scores typically drop compared to the original configuration, as indicated by negative EmD values. This performance dip is especially notable on the WTQ dataset, where even minor perturbations lead to reduced EM, and certain operations like column transpose and NT conditions cause pronounced declines in EM. Meanwhile, the VP metrics show that predictions shift more substantially under these transformations, signaling sensitivity in both large and small models to altered table structures. Although large models usually have higher baseline EM, they still exhibit performance degradation under perturbations. Particularly striking are conditions like NT—where the table is effectively removed—resulting in EM plummeting to near-zero and high VP, underscoring the models’ heavy reliance on the given tabular data. The results highlight the need for improved architectural and training strategies aimed at enhancing structural awareness and mitigating the negative effects of domain-specific perturbations, thereby making LLMs more robust and reliable in real-world table comprehension scenarios.
Operation WTQ Dataset TATQA Dataset SCITAB Dataset
EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - - 0.33 - - 0.20 - - 0.098 - - 0.132 - -
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11 - - - - - - - - - - - -
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15 - - - - - - - - - - - -
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23 - - - - - - - - - - - -
Table 4: The average EM, VP, and Emd of small and large LLMs on all perturbations operations

### v3

@@ Appendix G Additional Robustness Result
The table 1 summarizes the impact of various perturbation operations on the accuracy and robustness of larger and smaller LLMs models when performing TQA across three distinct datasets (WTQ, TAT-QA, and SCITAB). For each perturbation category, the EM scores typically drop compared to the original configuration, as indicated by negative EmD values. This performance dip is especially notable on the WTQ dataset, where even minor perturbations lead to reduced EM, and certain operations like column transpose and NT conditions cause pronounced declines in EM. Meanwhile, the VP metrics show that predictions shift more substantially under these transformations, signaling sensitivity in both large and small models to altered table structures. Although large models usually have higher baseline EM, they still exhibit performance degradation under perturbations. Particularly striking are conditions like NT, where the table is effectively removed, resulting in EM plummeting to near-zero and high VP, underscoring the models’ heavy reliance on the given tabular data. The results highlight the need for improved architectural and training strategies aimed at enhancing structural awareness and mitigating the negative effects of domain-specific perturbations, thereby making LLMs more robust and reliable in real-world table comprehension scenarios.
@@
G.1 Does newer models have better TQA abilities?
Fig. 15 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Does instruction tuning help TQA? The heat maps in Fig. 10 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. Figure 10(right) distinctly indicates that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
Appendix H Attention Matrix Measure
Sink rate: Gu et al. introduced the metric that measures the proportion of attention that “collapses" onto a specific token or a small set of tokens. A high sink rate indicates that the attention head focuses very narrowly, potentially ignoring other parts of the input.
$\text{Sink Rate}_{\epsilon}=\frac{1}{T}\sum_{j}\alpha_{1,j}^{(\ell,h)}>\epsilon,$
where $\alpha_{1,j}^{(\ell,h)}$ is the attention weight from the first token (token with index 1(<bos>)) to the j-th token in the $\ell$-th layer and $h$-th attention head.
Entropy: In the context of attention, entropy measures the diversity or distribution of attention across all tokens. High entropy means the attention is distributed more evenly, while low entropy suggests a concentrated attention pattern.
$H_{i}=-\sum_{j}\textbf{A}_{ij}\log(\textbf{A}_{ij})$
where $\textbf{A}_{ij}$ is the attention weight to the $j$-th token.
Frobenius Norm: The Frobenius norm of the attention matrix provides a measure of the overall magnitude or “strength” of attention. It is calculated as the square root of the sum of the squares of all elements in the matrix, indicating the total amount of attention being distributed.
$||\textbf{A}||_{F}=\sqrt{\sum_{i}\sum_{j}\textbf{A}_{ij}^{2}}$
Norm Difference: The difference between the attention matrices computed on the original (dev) and perturbed table(pert):
$\mathrm{NormDiff}\;=\;\bigl{\|}\mathbf{A}^{\text{dev}}\;-\;\mathbf{A}^{\text{pert}}\bigr{\|}_{F},$
where $\mathbf{A}^{dev}$ and $\mathbf{A}^{pert}$ are the attention matrices tokenized using the original table (‘dev”) and the perturbed table, respectively, and $\lVert\cdot\rVert_{F}$ is the Frobenius norm.
H.1 Cross-Model and Dataset Analysis of Perturbation Sensitivity
Figure 16: Spearman correlation between perturbation-induced attention-entropy change and EM degradation across 18 model–dataset pairs.
To evaluate the generality of our findings beyond the Llama3-8B-Instruct model on WTQ, we extend our correlation analysis between attention entropy change and EM performance degradation across a diverse set of model-dataset combinations. These include all the small models defined in 3,for which each models are tested on the WTQ, TATQA, and SCITAB datasets. For each configuration, we compute the Spearman correlation between the perturbation-induced changes in attention entropy and the corresponding drop in EM scores.
As shown in Figure 16, the positive correlation trend persists across nearly all settings. In particular, Instruct and chat variants of the models consistently show the strongest correlation values across datasets, reinforcing their heightened sensitivity to attention dispersion caused by structural perturbations. Notably, even non-instruct variants like Llama2-7B and Mistral-7B exhibit moderate to strong correlations, although the magnitude tends to be lower and more variable. We also find that SCITAB dataset results in most variation in the correlation, mainly due to the performance of the model, with the original table itself having a significantly low performance.
These consistent patterns across models and datasets confirm the robustness of our main claim: perturbation severity that induces greater shifts in attention entropy is reliably associated with a decline in model performance. This suggests that attention entropy change serves as a broadly applicable proxy for evaluating robustness in table-based QA models.
Moreover, this analysis demonstrates that the observed dynamics are not specific to any single dataset or model architecture. Instead, they reflect a more general representational vulnerability within current attention-based architectures when handling perturbed structured inputs.
H.2 Correlation within individual attention
Figure 17: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset, unsorted to depict actual correlation over each attention head.
Figure 18: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. The plot highlights a strong positive correlation for Layer 0, Head 10, indicating high sensitivity of this head to structural perturbations from 17
Figure 19: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. In contrast to the previous head, Layer 19, Head 23 shows minimal correlation from 17
We find that structural perturbations in tabular inputs affect the internal attention dynamics of the LLMs specifically, Llama-8B-Instruct model on WTQ, and how those dynamics relate to the performance. Figure 17 presents an unsorted Spearman correlation heatmap between per-head changes in attention entropy and the corresponding EM score differences, revealing substantial heterogeneity across all 32 heads and 32 layers. To unpack this variability, Figures 18 and 19 zoom in on the two extremal heads identified by the heatmap: Layer 00, Head $10$ exhibits a strong positive relationship, higher entropy shifts reliably predict larger drops in EM, whereas Layer 19, Head 23 shows significantly less correlation, indicating that its attention distribution remains largely impervious to structural pertubation. Together, these visualizations underscore that certain attention heads are sensitive for model robustness, while others maintain stable focus patterns even under severe perturbations.